# Think2Drive: Efficient Reinforcement Learning by Thinking in Latent World Model for Quasi-Realistic Autonomous Driving (in CARLA-v2)

### v1

@@ Abstract
Real-world autonomous driving (AD) especially urban driving involves many corner cases. The lately released AD simulator CARLA v2 adds 39 common events in the driving scene, and provide more quasi-realistic testbed compared to CARLA v1. It poses new challenge to the community and so far no literature has reported any success on the new scenarios in V2 as existing works mostly have to rely on specific rules for planning yet they cannot cover the more complex cases in CARLA v2. In this work, we take the initiative of directly training a planner and the hope is to handle the corner cases flexibly and effectively, which we believe is also the future of AD. To our best knowledge, we develop the first model-based RL method named Think2Drive for AD, with a world model to learn the transitions of the environment, and then it acts as a neural simulator to train the planner. This paradigm significantly boosts the training efficiency due to the low dimensional state space and parallel computing of tensors in the world model. As a result, Think2Drive is able to run in an expert-level proficiency in CARLA v2 within 3 days of training on a single A6000 GPU, and to our best knowledge, so far there is no reported success (100% route completion) on CARLA v2. We also propose CornerCase-Repository, a benchmark that supports the evaluation of driving models by scenarios. Additionally, we propose a new and balanced metric to evaluate the performance by route completion, infraction number, and scenario density, so that the driving score could give more information about the actual driving performance.
@@ 1 Introduction
Autonomous driving (AD) [18], especially urban driving, requires self-driving vehicles to handle regular tasks e.g. lane following and simple turns, as well as engage with traffic participants [21, 23, 22] and adapt to complex and dynamic traffic scenarios. With the increasing requirements from L2 level to L4, driving agents should be able to deal with high traffic complexity and intricate road topology.
@@
There have emerged simulators e.g. HighwayEnv [26] and CARLA v1 [8] for the development and validation of AD models. In these early simulators, the behaviors of environment agents are usually simple and the diversity and complexity of road conditions are usually limited, bearing a gap to real-world driving. Specifically, most of their tasks can be effectively addressed by rule-based approaches via basic skills like lane following, adherence to traffic signs, and collision avoidance, which is however well below the difficulty level of urban driving. For instance, in CARLA v1, the rule-based Autopilot with only hundreds of lines of code could achieve nearly perfect performance [20]. Thus methods developed for, or verified in these environments are possibly not able to deal with many common real-world traffic scenarios, which seriously limit their practical value.
@@
Accordingly, CARLA v2 [2] was released in November 2022, encompassing 39 real-world corner cases in addition to its v1 version of 10 cases. For instance, there are scenarios where the ego vehicle is on a two-way single-lane road and encounters a construction zone ahead. It requires the ego agent to invade the opposite lane when it is sufficiently clear, circumventing the construction area, and promptly merging back into the original lane afterward. In particular, corner cases, as their names suggest, are sparse in both the real world and the routes provided by CARLA v2, posing a long-tail problem for learning.
@@
Being aware of the difficulty of their benchmark, the CARLA team also provides several human demonstrations of completing these scenarios. Though humans could effortlessly navigate such scenarios, it is highly non-trivial to write into rules, not to mention adopting the popular imitation learning methods [19] with few samples as succeed in CARLA v1. Due to its much-increased difficulty, widely used rule-based experts like autopilot [20] and learning-based experts like Roach [38] (by model-free RL) both can not work in CARLA v2 at all. Up to date, there is no success reported on CARLA v2, one year after its release.
@@
In this paper, we aim to obtain the driving policy under such quasi-realistic AD simulations by learning, ambitiously with a model-based RL [15] approach which hopefully would enjoy two merits: data efficiency and flexibility against complex scenarios. It is well known that developing model-based RL can be nontrivial, depending on the specific domain area. It not only involves difficulties intrinsic to AD such as complex road conditions and highly interactive behaviors in long-tailed distributions, but also engineering problems of CARLA e.g. collecting massive samples efficiently from a cumbersome simulator. In contrast, to our best knowledge, existing RL methods for AD [3, 38] are all model-free [31, 27], which would inherently suffer from data inefficiency due to the complexity in CARLA v2.
@@
We model the environment’s transition function in AD using a world model [11, 14, 17] and employ it as a neural network simulator, to make the planner ‘think’ to drive (Think2Drive) in the learned latent space. In this way, the data efficiency could be significantly increased since the neural network could in parallel conduct hundreds of rollouts with a much faster iteration speed compared to the physical simulator i.e. CARLA. However, even with the state-of-the-art model-based methods [17], it is still highly non-trivial to adopt them for autonomous driving which has its unique characteristics as mentioned below compared to Atari or MineCraft as done in [17]. Specifically, we consider three major obstacles and our corresponding techniques in model-based RL for quasi-realistic AD.
@@
1) Policy degradation. There might exist contradictions among optimal policies of different scenarios. Consequently, the driving model will be easily trapped in the local optima. To mitigate this issue, we randomly re-initialize all weights of the planner in the middle of training while keeping the world model unchanged inspired by [28], allowing the planner to escape local optima for policy degradation preventing. As the world model can provide the planner with accurate and dense rewards after being well-trained, the reinitialized planner can better deal with the cold-start problem.
@@
2) Long-tail nature. As mentioned above, the long-tail nature of quasi-realistic AD tasks poses a significant challenge for the planner to handle all the corner cases. To this end, we implement an automated scenario generator that can generate scenarios based on road situations, thus providing the planner with abundant, scenario-dense data. We further design a termination-priority replay strategy, ensuring that the world model and planner prioritize exploration on long-tailed valuable states.
@@
Beyond these three major obstacles, the training of a model-based AD planner also encounters challenges such as initial startup difficulties, delayed learning signals, etc. We address them brick by brick, with a detailed discussion provided in Sec. 3.3. By developing all the above techniques, we manage to establish our model, Think2Drive, which has achieved the pioneering feat of successfully addressing all 39 quasi-realistic scenarios within 3 days of training on a single GPU A6000. Think2Drive aims to serve as an excellent planning module or teacher model for the development of advanced learning-based models for complex urban driving.
The highlights of the paper are as follows. 1) To our best knowledge, we are the first model-based RL approach for AD in literature that manages to handle quasi-realistic scenarios with tailored bricks including resetting technique, automated scenario generation, termination-priority replay strategy, steering cost function, etc.; 2) We propose a new and balanced metric to evaluate the performance by route completion and infraction number per kilometer; 3) Experimental results on CARLA V2 and the proposed CornerCaseRepository benchmark show the superiority of our approach.
@@ 2 Related Works
Model-based Reinforcement Learning. These methods explicitly utilize a world model to learn the transition of the environment and make the actor purely interact with the world model to improve data efficiency. PlaNet [14] proposes the recurrent state-space model (RSSM) to model both the deterministic and stochastic part of the environment followed by many later works [13, 16, 30, 17]. For instance, Dreamer [13], Dreamer2 [16], and Dreamer3 [17] progressively improve the performance of the world model based on RSSM. Notably, DreamerV3 achieves state-of-the-art performance across multiple tasks including Minecraft, DMLab, and Crafter, without the need of parameter tuning by employing techniques including symlog loss and free bits. Daydreamer [37] further extends the application of model-based approaches to physical robots.
We note that model-based RL is especially fit for AD since (i) the super data efficiency of the model-based method is the key to deal with the long-tailed issue of AD while the physical simulator is usually burdensome; (ii) the transition of AD scenes under rasterized BEV is relatively easy to learn compared to Atari or MineCraft, which means one would be able to train a highly accurate world model. However, it is non-trivial to directly adopt the model-based approach for self-driving because of some unique characteristics such as long-tail nature and reward contradictions.
Learning-based Agents in CARLA. In the traditional AD stack, the decision-making process is usually rule-based [33]. Though manual rules enable explainable control over the vehicle, it is difficult to deal with gradually appearing long-tailed corner cases. Learning-based agents aim to overcome the obstacle by computations. However, it is expensive and unsafe to directly test learning-based agents in the real world. The simulator CARLA [8] provides a basic environment. Among these works, [9] demonstrates the effectiveness of an end-to-end imitation learning (IL) baseline using a single-camera input. Additionally, approaches e.g. CIL [6] and CILRS [7] utilize branched action heads, selecting branches based on high-level directional commands. LBC [4] involves mapping camera images to waypoints, emulating a mid-to-mid imitation learning agent, which, in turn, utilized bird’s-eye view (BEV) data to generate future waypoints. [5] explores the utilization of BEV data as input for DDQN [34], TD3 [10], and SAC [12], with the added step of pre-training the image encoder on expert trajectories. [29] investigates the integration of IL with reinforcement learning. Roach [38] is the state-of-the-art model-free RL agent widely used as the expert model of recent end-to-end AD [35, 36, 25, 24], yet it fails to handle the CARLA v2 (Details in Sec. 4.4).
However, even Roach is not able to solve CARLA v2 and thus we propose Think2Drive, a model-based RL method, with tailored designs for quasi-realistic urban driving.
@@ 3.1 Problem Formulation with Model-based RL
We construct a planner model $\pi_{\eta}$ which outputs action $a_{t}$ based on current state $s_{t}$ and construct a world model $F$ to learn the transition of the driving scene so that the planner model could drive and be trained by “think” instead of directly interacting with the physical simulator. The iteration process of “think” is as follows: given an initial input $x_{t}$ at time-step $t$ sampled from the record, the world model encodes it as state $s_{t}$. Then, the planner generates $a_{t}$ based on $s_{t}$. Finally, the world model predicts the reward $r_{t}$, termination status $c_{t}$, and the future state $s_{t+1}$ with $s_{t}$ and $a_{t}$ as input. The overall pipeline is:
@@
By rollouting in the latent state space $s_{t}$ of the world model, the planner can think and learn efficiently, reducing the time cost of interacting with the heavy physical simulator.
@@ 3.2 World Model Learning and Planner Learning
We use DreamerV3 [17]’s structure and objective to train the world model and planner model.
World Model Learning. It has four components [17]:
RSSM{ Sequence model: ht=fθ(ht1,zt1,at1) Encoder: ztqθ(ztht,xt) Dynamics predictor: z^tpθ(z^tht) Reward predictor: r^tpθ(r^tht,zt)
 Termination predictor: c^tpθ(c^tht,zt) Decoder: x^tpθ(x^tht,zt)
@@
where RSSM is to provide an accurate transition function of the environment in latent space and perform efficient rollouts for the planner model. It decomposes the state representation $s_{t}$ into stochastic representation $z_{t}$ and deterministic hidden state $h_{t}$ based on Eq. 1 to better model the corresponding deterministic and stochastic aspects of the true transition function. The encoder first maps raw input $x_{t}$ to latent representation $z_{t}$, then the sequence model predicts the future hidden state $h_{t+1}$ based on the representation $z_{t}$, action $a_{t}$, and history hidden state $h_{t}$. The reward predictor forecasts the reward $r_{t}$ associated with the model state $s_{t}=(h_{t},z_{t})$ and the termination predictor predicts the termination flags $c_{t}\in\{0,1\}$, which both provide learning signals for the planner model. The decoder reconstructs inputs to ensure informative representation and generate interpretable images.
The world model’s training loss [17] consists of:
$\mathcal{L}_{\text{pred}}(\theta)\doteq$ $-\ln p_{\theta}\left(x_{t}\mid z_{t},h_{t}\right)-\ln p_{\theta}\left(r_{t}\mid z_{t},h_{t}\right)$ (3)
$-\ln p_{\theta}\left(c_{t}\mid z_{t},h_{t}\right)$
@@
As the horizon $T$ gets longer, the accumulated error exponentially grows. Thus, the expected return is clipped by $T=15$ and the left return is estimated by the critic: $R_{T:}=v(s_{T})$. We follow DreamerV3 by employing the trick of bucket-sorting rewards and the two-hot encoding to stably train the critic. Given two-hot encoded target $y_{t}=\operatorname{fz}(\text{twohot}(\text{symlog}(R_{t}^{\lambda}))))$, the cross-entropy loss is used to train the critic [17] :
@@
where the softmax distribution $p_{\psi}\left(\cdot\mid s_{t}\right)$ over equal split buckets is the output of the critic. The reward expectation term for actor training is normalized by moving statistics:
@@ 3.3 Challenges and Our Devised Bricks
After having the training paradigm determined, it is still not ready to solve the problem, as the autonomous driving scene has very different characteristics compared to Atari or MineCraft, e.g. policy degradation(Challenge 1), long-tail nature(Challenge 2, 3), and car heading stabilization(Challenge 4). These obstacles make it highly non-trivial to adopt MBRL for autonomous driving. To this end, we devise essential Bricks to address them one by one.
Challenge 1: As the training progresses, the agent may be trapped in the local optimal policy of easy scenarios. This issue arises from potential contradictions in optimal strategies required for different scenarios. For example, in scenario I where the front vehicle suddenly brakes and scenario II where the planner has to merge into high-speed traffic, the former requires the planner to keep a safe distance from the preceding vehicle, while the latter demands proactive engagement with the front vehicles. Since the former is much easier than the latter, the model can be easily trapped into the local optima of keeping safe distance.
@@
Challenge 2: The 39 scenarios are sparse in the released routes of CARLA v2. It brings a long-tail problem, which incurs skewed exploration over trivial states. Additionally, the released scenarios are coupled with specified waypoints. As a result, the scenarios happen in a few fixed locations, which limits data diversity.
Brick 2: We implement an automated scenario generator. Given a route, it can automatically split the route into multiple short routes and generate scenarios according to the road situation. As a result, the training process is able to acquire numerous shorter routes with dense scenarios. Besides, we build a benchmark CornerCaseRepository for evaluation which could estimate the detailed capabilities under each scenario while the official test routes are too long and thus their results are difficult to analyze.
@@
Brick 4: To address this, we incorporate a steering cost function into the training of our agent. This cost function has enabled our model to achieve stable navigation.
@@
Table 1: Driving performance and infraction of agents on the proposed SW benchmark. Mean and standard deviation are over 3 runs.
Input
Method
Driving
score
Weighted
DS
Route
completion
Infraction
penalty
Collision
pedestrians
Collision
vehicles
Collision
layout
Red light
Infraction
Stop sign
infraction
Agent
blocked
Privileged
Information
Roach [38] 57.5$\pm$9 54.8$\pm$0.5 96.4$\pm$1.1 0.59$\pm$0.28 0.85$\pm$0.56 8.42$\pm$4.65 0.85$\pm$0.51 0.56$\pm$0.45 0.49$\pm$0.44 0.78$\pm$0.31
@@
Raw Sensors
Think2Drive
+TCP [35]
36.40$\pm$12.23 29.6$\pm$0.2 85.88$\pm$8.26 0.41$\pm$0.32 0.46$\pm$0.32 9.92$\pm$5.12 6.75$\pm$3.08 3.27$\pm$1.64 5.03$\pm$3.82 6.18$\pm$4.62
@@
Scenario
Success Rate
Scenario
Success Rate
Scenario
Success Rate
Scenario
Success Rate
ParkingExit 0.89
Hazard
AtSidelane
0.75
Vinilla
Turn
0.99
Invading
Turn
0.90
Signalized
LeftTurn
0.95
Signalized
RightTurn
0.76
OppositeVehicle
TakingPriority
0.89
OppositeVehicle
RunningRedLight
0.85
Accident
0.81
Accident
TwoWays
0.61
Crossing
BicycleFlow
0.83
Highway
CutIn
1.0
Construction
0.84
Construction
TwoWays
0.72
Interurban
ActorFlow
0.83
InterurbanAdvanced
ActorFlow
0.8
Blocked
Intersection
0.80
Enter
ActorFlow
0.65
NonSignalized
RightTurn
0.75
NonSignalizedJunction
LeftTurnEnterFlow
0.67
MergerInto
SlowTraffic
0.67
MergerInto
SlowTrafficV2
0.87
Highway
Exit
0.83
NonSignalized
JunctionLeftTurn
0.79
SignalizedJunction
LeftTurnEnterFlow
0.86
Vehicle
TurningRoute
0.78
VehicleTurning
RoutePedestrian
0.75
Pedestrain
Crossing
0.91
YieldTo
EmergencyVehicle
0.92
Hard
Brake
1.0
Parking
CrossingPedestrian
0.98
Dynamic
ObjectCrossing
0.94
Vehicles
DooropenTwoWays
0.78
HazardAt
SideLaneTwoWays
0.92
Parked
Obstacle
0.90
ParkedObstacle
TwoWays
0.91
Static
CutIn
0.85
Parking
CutIn
0.90
ControlLoss
0.78

### v2

# Equal contributions  Correspondence (partly supported by NSFC 92370201)

@@ Abstract
Real-world autonomous driving (AD) like urban driving involves many corner cases. The lately released AD Benchmark CARLA Leaderboard v2 (a.k.a. CARLA v2) involves 39 new common events in the driving scene, providing a more quasi-realistic testbed compared to CARLA Leaderboard v1. It poses new challenges and so far no literature has reported any success on the new scenarios in V2. In this work, we take the initiative of directly training a neural planner and the hope is to handle the corner cases flexibly and effectively. To our best knowledge, we develop the first model-based RL method (named Think2Drive) for AD, with a compact latent world model to learn the transitions of the environment, and then it acts as a neural simulator to train the agent i.e. planner. It significantly boosts the training efficiency of RL thanks to the low dimensional state space and parallel computing of tensors in the latent world model. Think2Drive is able to run in an expert-level proficiency in CARLA v2 within 3 days of training on a single A6000 GPU, and to our best knowledge, so far there is no reported success (100% route completion) on CARLA v2. We also develop CornerCaseRepo, a benchmark that supports the evaluation of driving models by scenarios. We also propose a balanced metric to evaluate the performance by route completion, infraction number, and scenario density.
Keywords:
Autonomous driving Neural planner World model Model-based reinforcement learning CARLA v2 Think2Drive
We are getting to the point where there’s one last piece of the system that needs to be a neural net which is the planning and control function.
Elon Musk, 2023 Tesla Annual Shareholder Meeting
Vehicle control is the final piece of the Tesla FSD AI puzzle. That will drop >300k lines of C+++++ + control code by  2 orders of magnitude. It is training as I write this.
Elon Musk, Twitter August, 2023
@@ 1 Introduction
Autonomous driving (AD) [29, 18, 44], especially urban driving, requires the vehicles to engage with dense and diverse traffic participants [24, 22, 25] and adapt to complex and dynamic traffic scenarios. Traditional manually-crafted rule-based planning methods struggle to handle these scenarios due to their reliance on exhaustive coverage of all cases, which is nearly impossible for long-tail scenarios [27, 31]. Additionally, ensuring compatibility between new and existing rules becomes increasingly challenging as the decision tree expands [2]. As a result, there is a trend that to adopt neural planner, offering hope for scaling up with data and computation to achieve full driving autonomy.
@@
There have emerged benchmarks for the development and validation of planning methods in AD, e.g. HighwayEnv [28] and CARLA Leaderboard v1 (a.k.a. CARLA v1) [8]. However, in these pioneering benchmarks, the behaviors of environment agents are usually simple and the diversity and complexity of road conditions are often limited, bearing a gap to real-world driving. In fact, most of their tasks can be effectively addressed by rule-based approaches [6] via basic skills like lane following, adherence to traffic signs, and collision avoidance, which is however well below the difficulty level of real-world urban driving. For instance, in CARLA v1, the rule-based Autopilot with only hundreds of lines of code could achieve nearly perfect performance [20]. Thus, methods developed for or verified in these environments are possibly unable to handle many common real-world traffic scenarios, which significantly limits their practical value.
@@
A quasi-realistic benchmark, CARLA Leaderboard v2 (a.k.a. CARLA v2) [3] was released in November 2022, encompassing 39 real-world corner cases in addition to its v1 version with 10 cases. For instance, there are scenarios where the ego vehicle is on a two-way single-lane road and encounters a construction zone ahead. It requires the ego agent to invade the opposite lane when it is sufficiently clear, circumventing the construction area, and promptly merging back into the original lane afterward. In particular, corner cases, as their names suggest, are sparse in both the real world and the routes provided by CARLA v2, posing a long-tail problem for learning.
@@
Being aware of the difficulty of the new benchmark, the CARLA team also provides several human demonstrations of completing these scenarios. Though humans could effortlessly navigate such scenarios, it is highly non-trivial to write into rules, not to mention adopting the popular imitation learning methods [19] with few samples, as succeed in CARLA v1. Due to the much-increased difficulty, widely used rule-based experts like autopilot [20] and learning-based experts like Roach [45] (by model-free RL) both can not work in CARLA v2 at all. Up to date, there is no success reported on CARLA v2, one year after its release.
@@
In this paper, we aim to obtain the driving policy under such a quasi-realistic AD benchmark by learning, ambitiously with a model-based RL [13] approach which hopefully would enjoy two merits: data efficiency and flexibility against complex scenarios. Note that developing model-based RL can be nontrivial, w.r.t. the specific domain. It not only involves difficulties intrinsic to AD such as complex road conditions and highly interactive behaviors in long-tailed distributions, but also engineering problems of CARLA e.g. collecting massive samples efficiently from a cumbersome simulator. In contrast, to our best knowledge, existing RL methods for AD [4, 45] are mostly model-free [36, 32], which would inherently suffer from data inefficiency in CARLA v2 due to its complexity. There are also a few model-based RL methods applied to AD [17, 7] in simple benchmarks, but far from solving CARLA v2.
Specifically, we model the environment’s transition function in AD using a world model [10, 14, 16] and employ it as a neural network simulator, to make the planner ‘think’ to drive (Think2Drive) in the learned latent space.
@@
In this way, the data efficiency could be significantly increased since the neural network could in parallel conduct hundreds of rollouts with a much faster iteration speed compared to the physical simulator i.e. CARLA. However, even with the state-of-the-art model-based methods [16], it is still highly non-trivial to adopt them for AD which has its unique characteristics as mentioned below compared to Atari or MineCraft as done in [16]. Specifically, we consider three major obstacles in model-based RL for quasi-realistic AD.
@@
1) Policy degradation. There might exist contradictions among optimal policies of different scenarios. For instance, for scenario I where the front vehicle suddenly brakes, and scenario II where the planner has to merge into high-speed traffic, the former requires the planner to keep a safe distance from the preceding vehicle, while the latter demands proactive engagement with the front vehicles. Consequently, the driving model will be easily trapped in the local optima. To mitigate this issue, we randomly re-initialize all weights of the planner in the middle of training while keeping the world model unchanged inspired by [33], allowing the planner to escape local optima for policy degradation preventing. As the world model can provide the planner with accurate and dense rewards, the reinitialized planner can better deal with the cold-start problem.
@@
2) Long-tail nature. As mentioned above, the long-tail nature of AD tasks poses a significant challenge for the planner to handle all the corner cases. We implement an automated scenario generator that can generate scenarios based on road situations, thus providing the planner with abundant, scenario-dense data. We further design a termination-priority replay strategy, ensuring that the world model and planner prioritize exploration on long-tailed valuable states.
@@
Beyond these three major obstacles, the training of a model-based AD planner also encounters challenges such as initial running difficulties, delayed learning signals, etc. We address them brick by brick, with a detailed discussion provided in Sec. 3.3. By developing all the above techniques, we manage to establish our model, Think2Drive, which has achieved the pioneering feat of successfully addressing all 39 quasi-realistic scenarios within 3 days of training on a single GPU A6000. Think2Drive can also serve as a planning module or teacher model for learning-based driving models.
The highlights of the paper are as follows. 1) To our best knowledge, it is the first model-based RL approach for AD (i.e. neural planner) in literature that manages to handle quasi-realistic scenarios with techniques like resetting technique, automated scenario generation, termination-priority replay strategy, steering cost function, etc. 2) We propose a new and balanced metric to evaluate the performance by route completion, infraction number and scenario density. 3) Experimental results on CARLA V2 and the proposed CornerCaseRepo benchmark show the superiority of our approach. A demo of our proficient planner on CARLA V2 test routes is available at https://thinklab-sjtu.github.io/CornerCaseRepo/
@@ 2 Related Works
Model-based Reinforcement Learning. Model-based reinforcement learning explicitly utilizes a world model to learn the transition of the environment and make the actor purely interact with the world model to improve data efficiency. PlaNet [14] proposes the recurrent state-space model (RSSM) to model both the deterministic and stochastic part of the environment followed by many later works [12, 15, 35, 16]. For instance, Dreamer [12], Dreamer2 [15], and Dreamer3 [16] progressively improve the performance of the world model based on RSSM. Notably, DreamerV3 achieves state-of-the-art performance across multiple tasks including Minecraft and Crafter, without the need of parameter tuning by employing techniques including symlog loss and free bits. Daydreamer [43] further extends the application of model-based approaches to physical robots.
We note that model-based RL is especially fit for AD since 1) the super data efficiency of the model-based method could be the key to deal with the long-tailed issue of AD while the physical simulator is usually burdensome; 2) the transition of AD scenes under rasterized BEV is relatively easy to learn compared to Atari or MineCraft, which means one would be able to train an accurate world model.
Reinforcement Learning-based Agents in CARLA. Reinforcement learning is an important technique to obtain planning agents in CARLA, which could serve as expert models. [5] explores the utilization of BEV data as input for DDQN [39], TD3 [9], and SAC [11], with the added step of pre-training the image encoder on expert trajectories. [34] investigates the integration of IL with reinforcement learning. MaRLn [38] uses raw sensor inputs to train a reinforcement learning (RL) agent, failing to achieve good performance in Leaderboard v1. Roach [45] is the state-of-the-art model-free RL agent widely used as the expert model of recent end-to-end AD [42, 41, 26, 23], yet it fails to handle the CARLA v2 (details in Sec. 4.4). There are also a few model-based RL methods applied to AD [17, 7] in simple benchmarks, but the difficulty is far from CARLA v2. Notably, after the publishing of Think2Drive, PDM-lite [21], a recently released rule-based planner (after our arxiv version [30]), is able to solve all scenarios in CARLA Leaderboard v2 as well. However, they adopt different hyper-parameters for different scenarios, which requires heavy manual labor.
@@ 3.1 Problem Formulation with Model-based RL
We construct a planner model $\pi_{\eta}$ which outputs action $a_{t}$ based on current state $s_{t}$ and construct a world model $F$ to learn the transition of the driving scene so that the planner model could drive and be trained by “think" instead of directly interacting with the physical simulator. The iteration process of “think" is as follows: given an initial input $x_{t}$ at time-step $t$ sampled from the record, the world model encodes it as state $s_{t}$. Then, the planner generates $a_{t}$ based on $s_{t}$. Finally, the world model predicts the reward $r_{t}$, termination status $c_{t}$, and the future state $s_{t+1}$ with $s_{t}$ and $a_{t}$ as input. The overall pipeline is:
@@
By rollouting in the latent state space $s_{t}$ of the world model, the planner can think and learn efficiently, without interacting with the heavy physical simulator.
@@ 3.2 World Model Learning and Planner Learning
We use DreamerV3 [16]’s structure and objective to train the world model and planner model. Note that our main novelty lies in the first successful adoption of latent world model to AD.
World Model Learning. It has four components in line with [16]:
@@
[…]
where RSSM is to provide an accurate transition function of the environment in latent space and perform efficient rollouts for the planner model. It decomposes the state representation $s_{t}$ into stochastic representation $z_{t}$ and deterministic hidden state $h_{t}$ based on Eq. 1 to better model the corresponding deterministic and stochastic aspects of the true transition function. The encoder first maps raw input $x_{t}$ to latent representation $z_{t}$, then the sequence model predicts the future hidden state $h_{t+1}$ based on the representation $z_{t}$, action $a_{t}$, and history hidden state $h_{t}$. The reward predictor forecasts the reward $r_{t}$ associated with the model state $s_{t}=(h_{t},z_{t})$ and the termination predictor predicts the termination flags $c_{t}\in\{0,1\}$, which both provide learning signals for the planner model. The decoder reconstructs inputs to ensure informative representation and generate interpretable images. Specifically, the world model’s training loss [16] consists of:
$\mathcal{L}_{\text{pred}}(\theta)\doteq$ $-\ln p_{\theta}\left(x_{t}\mid z_{t},h_{t}\right)-\ln p_{\theta}\left(r_{t}\mid z_{t},h_{t}\right)-\ln p_{\theta}\left(c_{t}\mid z_{t},h_{t}\right)$ (3)
@@
To handle the accumulated error over the horizon $T$, the expected return is clipped by $T=15$ and the left return is estimated by the critic: $R_{T:}=v(s_{T})$. We follow DreamerV3 by employing the bucket-sorting rewards and two-hot encoding to stably train the critic. Given the two-hot encoded target $y_{t}=\operatorname{fz}(\text{twohot}(\text{symlog}(R_{t}^{\lambda}))))$, the cross-entropy loss is used to train the critic [16]:
@@
where the softmax distribution $p_{\psi}\left(\cdot\mid s_{t}\right)$ over equal split buckets is the output of the critic. The reward expectation term for actor training is normalized by moving statistics [15, 40]:
@@ 3.3 Challenges and Our Devised Bricks
After having the training paradigm determined, it is still not ready to solve the problem, as the autonomous driving scene has very different characteristics compared to Atari or MineCraft, e.g. policy degradation(Challenge 1), long-tail nature(Challenge 2,3), car heading stabilization(Challenge 4), and some other obstacles(Challenge 5,6,7). These obstacles make it highly non-trivial to adopt MBRL for AD. We devise essential Bricks to address them one by one.
Challenge 1: During training, the agent may be trapped in the local optimal policy of easy scenarios. This issue arises from potential contradictions in optimal strategies required for different scenarios. For example, in scenario I where the front vehicle suddenly brakes and scenario II where the planner has to merge into high-speed traffic, the former requires the planner to keep a safe distance from the preceding vehicle, while the latter demands proactive engagement with the front vehicles. Since the former is much easier than the latter, the model can be easily trapped into the local optima of keeping safe distance.
@@
Challenge 2: The 39 scenarios are sparse in the released routes of CARLA v2. It brings a long-tail problem, incurring skewed exploration over trivial states. Also, the released scenarios are coupled with specified waypoints. As a result, the scenarios happen in a few fixed locations with limited diversity.
Brick 2: We implement an automated scenario generator. Given a route, it can automatically split the route into multiple short routes and generate scenarios according to the road situation. As a result, the training process is able to acquire numerous shorter routes with dense scenarios. Besides, we build a benchmark CornerCaseRepo for evaluation which could estimate the detailed capabilities under each scenario while the official test routes are too long and thus their results are difficult to analyze.
@@
Brick 4: We incorporate a steering cost function into the training of our agent. This cost function has enabled our model to achieve stable navigation.
@@
Table 1: Driving performance and infraction of agents on the proposed CornerCaseRepo benchmark. Mean and standard deviation are over 3 runs.
Input Method Driving score Weighted DS Route completion Infraction penalty Collision pedestrians Collision vehicles Collision layout Red light Infraction Stop sign infraction Agent blocked
Privileged Information Roach [45] 57.5$\pm$9 54.8$\pm$0.5 96.4$\pm$1.1 0.59$\pm$0.28 0.85$\pm$0.56 8.42$\pm$4.65 0.85$\pm$0.51 0.56$\pm$0.45 0.49$\pm$0.44 0.78$\pm$0.31
@@
Raw Sensors Think2Drive +TCP [42] 36.40$\pm$12.23 29.6$\pm$0.2 85.88$\pm$8.26 0.41$\pm$0.32 0.46$\pm$0.32 9.92$\pm$5.12 6.75$\pm$3.08 3.27$\pm$1.64 5.03$\pm$3.82 6.18$\pm$4.62
@@
Scenario Success Rate Scenario Success Rate Scenario Success Rate Scenario Success Rate
ParkingExit 0.89 Hazard AtSidelane 0.75 Vinilla Turn 0.99 Invading Turn 0.90
Signalized LeftTurn 0.95 Signalized RightTurn 0.76 OppositeVehicle TakingPriority 0.89 OppositeVehicle RunningRedLight 0.85
Accident 0.81 Accident TwoWays 0.61 Crossing BicycleFlow 0.83 Highway CutIn 1.0
Construction 0.84 Construction TwoWays 0.72 Interurban ActorFlow 0.83 InterurbanAdvanced ActorFlow 0.8
Blocked Intersection 0.80 Enter ActorFlow 0.65 NonSignalized RightTurn 0.75 NonSignalizedJunction LeftTurnEnterFlow 0.67
MergerInto SlowTraffic 0.67 MergerInto SlowTrafficV2 0.87 Highway Exit 0.83 NonSignalized JunctionLeftTurn 0.79
SignalizedJunction LeftTurnEnterFlow 0.86 Vehicle TurningRoute 0.78 VehicleTurning RoutePedestrian 0.75 Pedestrain Crossing 0.91
YieldTo EmergencyVehicle 0.92 Hard Brake 1.0 Parking CrossingPedestrian 0.98 Dynamic ObjectCrossing 0.94
Vehicles DooropenTwoWays 0.78 HazardAt SideLaneTwoWays 0.92 Parked Obstacle 0.90 ParkedObstacle TwoWays 0.91
Static CutIn 0.85 Parking CutIn 0.90 ControlLoss 0.78